i am 600 miles away from home traveling at 65 mph and i need an equation to find out when ill be 300 miles away from home. is y=65x+600 a good equation?

Answer by Alan3354(69443)

i am 600 miles away from home traveling at 65 mph and i need an equation to find out when ill be 300 miles away from home. is y=65x+600 a good equation?
--------------
You didn't state which way you were traveling.
If you're moving away, it's good.
Notice that y increases with x.
-----------------
Asking when you'll be 300 miles away implies you're traveling towards home.
--> d = 600 - 65x
300 = 600 - 65x
65x = 300
x = 300/65 = 60/13 hours